OpenStudy (anonymous):

If the Federal Reserve decreases the reserve rate from 7% to 5%, how does this affect the amount of money that would result because of fractional-reserve banking from an initial deposit into a bank of $30,000?

OpenStudy (tkhunny):

30000*0.07 = 30000*0.05 =